# Learning Lens Blur Fields

### v1

@@ Abstract
Optical blur is an inherent property of any lens system and is challenging to model in modern cameras because of their complex optical elements. To tackle this challenge, we introduce a high-dimensional neural representation of blur—the lens blur field—and a practical method for acquiring it. The lens blur field is a multilayer perceptron (MLP) designed to (1) accurately capture variations of the lens 2D point spread function over image plane location, focus setting and, optionally, depth and (2) represent these variations parametrically as a single, sensor-specific function. The representation models the combined effects of defocus, diffraction, aberration, and accounts for sensor features such as pixel color filters and pixel-specific micro-lenses. To learn the real-world blur field of a given device, we formulate a generalized non-blind deconvolution problem that directly optimizes the MLP weights using a small set of focal stacks as the only input. We also provide a first-of-its-kind dataset of 5D blur fields—for smartphone cameras, camera bodies equipped with a variety of lenses, etc. Lastly, we show that acquired 5D blur fields are expressive and accurate enough to reveal, for the first time, differences in optical behavior of smartphone devices of the same make and model.
@@ I Introduction
Despite the potentially wide-ranging impact of such models, existing approaches for estimating optical blur in real-world camera systems trade-off representational accuracy against scalability and ease of acquisition. To address this major limitation, we propose a novel method for blur acquisition that enables highly accurate modeling of smartphone and interchangeable-lens camera systems without sacrificing scalability across many potential dimensions (2D pixel location, depth, focus setting, wavelength, etc.) or ease of use. A key property of our approach is that it outputs a model that is continuous, making it possible to represent both large scale and minute blur variations across all those dimensions. Although our specific focus in this paper is conventional cameras for which a thin-lens model can serve as a (very crude) approximation, generalization to systems employing coded optics [13, 14, 15, 16, 17] may be possible as well.
@@ I-A Background and Related Work
Non-parametric methods sample the space of a camera’s PSFs by acquiring a sparse grid of 2D PSFs over the sensor plane [31, 32, 22, 33] and optionally over focus settings [31] and depth [32]. These approaches are fundamentally limited by a density-accuracy tradeoff because they solve an independent, local blind [31, 24, 34, 35, 36, 37] or non-blind [38, 39, 40, 41, 42] deconvolution problem for each PSF, which becomes degenerate as the image patches used for deconvolution become smaller and/or the PSFs spatial extent becomes larger [32, 43]. As a result, these methods sample PSFs sparsely, limiting their ability to model PSFs that vary rapidly along one or more dimensions [44]. Although in theory such limitations can be side-stepped by imaging an actual point light source, such an approach comes with its own major limitations, including light efficiency and the inability to create an ideal point source because of diffraction [2, 45].
Optical simulation systems [46, 47, 48, 49, 50], on the other hand, assume that a physical model of the complete camera-plus-lens system is known precisely and that the corresponding real-world devices conform to it exactly. In many cases, however, precise models of widely-used compound lens systems may be proprietary or unknown [51], and their precision cannot be taken for granted due to device age, mechanical impacts from daily use, manufacturing variabilities, etc. While many deep learning techniques tackle the blind [52, 53, 54, 55, 56] or non-blind [57, 58, 59] deconvolution problems, none of these methods apply to the high-dimensional cases we consider. Moreover, these methods rely on a discrete PSF representation instead of an MLP and thus suffer from the curse of dimensionality as PSF dimensions increase. Consequently, the recovery of accurate, densely-sampled, high-dimensional PSFs for commodity cameras has remained beyond the reach of the broader computational photography community due to these technical obstacles.
@@ I-B Contributions
In this paper, we seek to fill this gap by introducing a novel neural representation of blur—the lens blur field—along with a practical method of acquiring it. The lens blur field represents a camera’s PSF as a multi-layer perceptron (MLP) that is designed to capture blur variations across sensor position, focus setting and, optionally, depth. This representation is particularly suitable for PSF acquisition because it yields a continuous representation of the PSF across multiple dimensions instead of sparse discrete samples; it can be optimized by leveraging the advanced MLP training schemes already available; and it can accurately capture difficult-to-model effects such as defocus, chromatic aberration, and higher-order aberrations with relatively few parameters.
@@
Our work provides new insights and enables new applications. For example we show that PSFs vary considerably between smartphone models, including the iPhone 12 Pro and iPhone 14 Pro; that dual-pixel auto-focus systems have PSFs that depart significantly from the analytical models used hitherto; and that lens blur fields are reliable and expressive enough to reveal subtle differences in the optical behavior of smartphone devices of the same make and model, yielding a device-specific “blur signature”. We also show that lens blur fields can be used as space-efficient computational proxies of device optics, allowing us to apply high-dimensional blurs to render images with device-specific PSFs (Figure 8). Taken together, we provide new ways to both model and simulate the performance of common imaging systems, including smartphone cameras used by billions worldwide.
Figure 2: Overview of learning lens blur fields. (A) Our approach learns a blur field by solving a non-blind deconvolution problem between blur-free calibration patterns $\hat{I}$ and blurry images $I$ of the patterns. Blur-free images are computed by applying perspective projection, lens distortion, and radiometric compensation to the albedo $A$ at the scene plane (top left). (B) An MLP is used to represent the lens blur field, which parameterizes the PSF over sensor position $\mathbf{x}$, target distance $d_{i}$, focus distance $f_{i}$, and a displacement $\mathbf{u}$ from $\mathbf{x}$. We train the MLP on a dataset of blur-free and captured focal stack images of calibration patterns and optimize MLP parameters to minimize the error between predicted and captured blurry image patches.
@@ II-A Lens Blur Field Representation
Consider an incoherent point light source $\mathbf{p}\in\mathbb{R}^{3}$ at distance $d$ from the front pupil plane of a lens (see Figure 2). Let $\mathbf{\tilde{x}}\in\mathbb{R}^{2}$ be the perspective projection of point $\mathbf{p}$ onto the camera’s sensor plane and let $\mathbf{x}\in\mathbb{R}^{2}$ be the point’s actual projection after accounting for non-linear geometric distortions caused by the lens. Since a camera’s focus setting $f$ may affect the overall magnification of captured images [60] and since dual-pixel AF sensors induce disparity between left-looking and right-looking pixels, in general $\mathbf{\tilde{x}}$ and $\mathbf{x}$ depend on both $f$ and the pixel type $c$. In the following, we express the one-to-one relation between geometrically-distorted points on the sensor plane and point sources at distance $d$ using
@@ II-B Continuous Image Formation
$\hat{I}^{(c)}(\mathbf{x},d,f)~{}=~{}A^{(c)}(\mathbf{p})~{}E^{(c)}(\mathbf{p},f)$ (2)
@@
As illustrated in Figure 2, we optimize the parameters of the blur field MLP by minimizing a photometric loss between a set of captured blurry images of known 2D patterns and corresponding rendered images estimated using Equation 3.
@@
The outer summation is performed over all pixel channels $c$, all sensor points $\mathbf{x}$, focus and distance settings $f_{i}$ and $d_{i}$, $1\leq i\leq N_{I}$, and for each of $N_{T}$ known patterns ($1\leq t\leq N_{T}$). The inner summation is an approximation of the integral formulation of Equation 3 with a discrete sum over $\mathbf{u}$. We consider two instances of the training procedure depending on whether or not the distance $d_{i}$ is the same for all input images. In the former case $d$ is fixed, yielding a 5D blur field; otherwise we optimize for a 6D blur field.
@@
We train on raw captured images, with the summation over $\mathbf{x}$ in Section II-B taken only over pixels for which we have image measurements for a given channel.
For example, for a sensor with a Bayer color filter array, we supervise the red, green, and blue channels only on coordinates $\mathbf{x}$ that are explicitly measured.
The MLP uses seven fully-connected layers of 512 channels each with ReLU activation functions. The final layer has a sigmoid activation to satisfy the non-negativity constraint. Unlike recent work employing MLPs, we do not apply positional encoding to the MLP’s input coordinates [63] as we find it unnecessary for accurate PSF acquisition. Blur fields are trained for between $4\times 10^{6}$ and $10\times 10^{6}$ iterations with the Adam optimizer ($\beta_{1}=0.5$, $\beta_{2}=0.999$) on a single NVIDIA A6000 using the tiny-cuda-nn framework [64]. Details on other hyperparameters (learning rate, exponential decay, etc.) are included in the supplement.
@@
During training, we set one batch element to comprise a 2D PSF obtained by first randomly sampling a focus setting $f_{i}$ and a point $\mathbf{x}$ at a pixel center. Then we sample a grid of points in the two dimensions of $\mathbf{u}$ within the patch around $\mathbf{x}$ and query the MLP at these points to retrieve a 2D PSF. Section II-B is computed using the output of the convolution of the 2D PSF with $\hat{I}$ within the area of the patch. In our experiments, we use a batch size of one and allow 2D PSFs up to $120\times 120$ samples in size. The size of a patch $\mathcal{P}$ is set to be 1.5 times the size of the 2D PSF along each dimension, and the convolution is performed within the valid region of the patch to avoid boundary artifacts.
@@ III Image Acquisition Procedure
We describe the procedure to acquire the blur-free and captured blurry images of the calibration patterns used to optimize the lens blur field. The steps of the procedure are outlined in Figure 2 (A) and consist of first computing the blur-free calibration image by applying (1) perspective projection via homography, (2) lens distortion, and (3) radiometric compensation to the scene-plane albedo $A$. We also capture the blurry images of the calibration patterns displayed on a monitor using a procedure outlined in this .
@@ III-A Homography Estimation
The perspective projection of the scene-plane albedo $A(\mathbf{p})$ onto the sensor plane (i.e., $\mathbf{p}\rightarrow\mathbf{\tilde{x}}$—see Figure 2) is computed via a homography $H(\mathbf{\tilde{x}},d_{i},c)$. To estimate the homography, we require a pattern that enables robust point localization with sub-pixel accuracy, even in the presence of large defocus. To this end, we use two dot-grid calibration patterns consisting of white circles on a black background and black circles on a white background (shown, right). We compute the albedo dot-grid images and capture in-focus dot-grid images for which we also remove radial distortion using the procedure described in the next sub-. Following Scharstein and Szeliski [66], we binarize the dot-grid images and detect the pixel coordinates of the dot centers in each image. Then we refine the dot center locations in both images using a center-of-mass calculation on the dots in the original albedo and captured images. Finally, we compute the homography using RANSAC [67] between the dot center point correspondences in the images. Applying the homography to the scene-plane albedo as $A(H^{-1}(\mathbf{\tilde{x}},d_{i},c))$ gives the sensor-plane albedo.
@@ III-B Lens Distortion Estimation
To estimate the lens distortion we use the sixth-order Brown-Conrady model of radial distortion implemented in OpenCV’s calibrateCamera function [68] and apply it to the dot-grid center positions of the captured in-focus image described in the homography estimation step. We apply this model of radial distortion $D(\mathbf{\tilde{x}},c)$ to all scaled sensor-plane albedo images over $f_{i}$ and $d_{i}$. The composition of homography $H$, scale transformation $S$, and radial distortion $D$ define the invproj operator introduced in Equation 1, which maps points $\mathbf{x}$ on the sensor plane to points $\mathbf{p}$ on the scene plane.
@@ III-C Radiometric Compensation
By combining Equation 4 and its analogue for $I_{\text{max}}$ which follows trivially, the blur-free image $\hat{I}$ can now be expressed as an affine combination of $I_{\text{min}}$ and $I_{\text{max}}$.
@@ III-D Image Capture
Finally, we describe the procedure for capturing focal stack images displayed on a monitor. We use a 32-inch 4K LED monitor to display the calibration patterns for both smartphone and SLR captures. We mount each camera on a tripod and position it fronto-parallel to the monitor. Depending on the distance $d$ to the scene plane, the image of the calibration pattern may span the entire sensor plane or just a portion of it. In the former case, a single focal stack of images is captured for lens distortion estimation and radiometric compensation. In the latter, the camera and/or monitor are physically moved to maintain $d$ while also adjusting the patterns to project to different parts of the sensor plane (see result in Section IV). For smartphone experiments, focal stacks are automatically captured using software, and we sample a range of focus positions denoted as $N_{I}\in\{7,24\}$, which depends on hardware and device limitations. Finally, we set the ISO to the lowest possible value and capture bursts with exposure times of 100 ms (for iPhones) or 250 ms (for Pixel and SLR) to avoid capturing screen refresh, PWM, and flicker. The supplement contains additional details on our acquisition procedure for all devices in our dataset.
Figure 3: Simulated recovery of spatially invariant dual-pixel blur kernels created with the parametric models proposed by Punnappurath et al. [10]. Each sub-panel gives a side-by-side comparison of the ground truth (left), our recovered PSF (middle left), the result of Joshi et al. [31] (middle right), and Mannan & Langer [32] (right). Root-mean-square error (RMSE) is estimated on the inset regions shown in Figure 3 to mitigate bias from large regions of zero-valued pixels. The values are shown in the bottom left corner of each plot. Supplement S5 contains more extensive comparisons and cross-sectional plots.

### v2

@@ Abstract
Optical blur is an inherent property of any lens system and is challenging to model in modern cameras because of their complex optical elements. To tackle this challenge, we introduce a high-dimensional neural representation of blur—the lens blur field—and a practical method for acquiring it. The lens blur field is a multilayer perceptron (MLP) designed to (1) accurately capture variations of the lens 2D point spread function over image plane location, focus setting and, optionally, depth and (2) represent these variations parametrically as a single, sensor-specific function. The representation models the combined effects of defocus, diffraction, aberration, and accounts for sensor features such as pixel color filters and pixel-specific micro-lenses. To learn the real-world blur field of a given device, we formulate a generalized non-blind deconvolution problem that directly optimizes the MLP weights using a small set of focal stacks as the only input. We also provide a first-of-its-kind dataset of 5D blur fields—for smartphone cameras, camera bodies equipped with a variety of lenses, etc. Lastly, we show that acquired 5D blur fields are expressive and accurate enough to reveal, for the first time, differences in optical behavior of smartphone devices of the same make and model. Code and data can be found at blur-fields.github.io.
@@ I Introduction
Despite the potentially wide-ranging impact of such models, existing approaches for estimating optical blur in real-world camera systems trade-off representational accuracy against ease of acquisition. Accurately representing optical blur across multiple dimensions requires densely sampling the parameter space, which increases acquisition time and estimation complexity. To address this major limitation, we propose a novel method for blur acquisition that enables highly accurate modeling of smartphone and interchangeable-lens camera systems without sacrificing scalability across many potential dimensions (2D pixel location, depth, focus setting, wavelength, etc.) or ease of use. A key property of our approach is that it outputs a model that is continuous, making it possible to represent both large scale and minute blur variations across all those dimensions. Although our specific focus in this paper is conventional cameras for which a thin-lens model can serve as a (very crude) approximation, generalization to systems employing coded optics [13, 14, 15, 16, 17] may be possible as well.
@@ I-A Background and Related Work
Non-parametric methods sample the space of a camera’s PSFs by acquiring a sparse grid of 2D PSFs over the sensor plane [31, 32, 22, 33] and optionally over focus settings [31] and depth [32]. These approaches are fundamentally limited by a density-accuracy tradeoff because they solve an independent, local blind [31, 24, 34, 35, 36, 37] or non-blind [38, 39, 40, 41, 42] deconvolution problem for each PSF, which becomes degenerate as the image patches used for deconvolution become smaller and/or the PSFs spatial extent becomes larger [32, 43]. As a result, these methods sample PSFs sparsely, limiting their ability to model PSFs that vary rapidly along one or more dimensions [44]. Although in theory these limitations can be side-stepped by adhering to the PSF definition and imaging a point light source to capture the PSF directly, in practice it is difficult to construct an ideal point light source [45], i.e., a source with a perfectly circular shape that projects to a region smaller than a pixel and has uniform radiant intensity. Moreover, sources with the micron-scale sizes required for the tiny pixels in modern smartphones are subject to diffraction [46, 47] and have very low light efficiency, making it even more difficult to treat their images as that of a “ground truth PSF”. For all these reasons, accurate estimation of spatially varying PSFs has long relied on deconvolution of known patterns [31] as opposed to direct imaging.
Optical simulation systems [48, 49, 50, 51, 52], on the other hand, assume that a physical model of the complete camera-plus-lens system is known precisely and that the corresponding real-world devices conform to it exactly. In many cases, however, precise models of widely-used compound lens systems may be proprietary or unknown [53], and their precision cannot be taken for granted due to device age, mechanical impacts from daily use, manufacturing variabilities, etc. While many deep learning techniques tackle the blind [54, 55, 56, 57, 58, 59] or non-blind [60, 61, 62, 63, 64] deconvolution problems, none of these methods apply to the high-dimensional, spatially-varying cases we consider. Moreover, these methods rely on a discrete PSF representation instead of an MLP and thus suffer from the curse of dimensionality as PSF dimensions increase. Consequently, the recovery of accurate, densely-sampled, high-dimensional PSFs for commodity cameras has remained beyond the reach of the broader computational photography community due to these technical obstacles.
@@ I-B Contributions
In this paper, we seek to fill this gap by introducing a novel neural representation of blur—the lens blur field—along with a practical method of acquiring it. The lens blur field represents a camera’s PSF as a multi-layer perceptron (MLP) that is designed to capture blur variations across sensor position, focus setting and, optionally, depth. This representation is particularly suitable for PSF acquisition because it yields a continuous representation of the PSF across multiple dimensions instead of sparse discrete samples; it can be optimized by leveraging the advanced MLP training schemes already available; it can learn high-dimensional signals more accurately while using only a fraction of the parameters of non-parametric approaches [65, 66]; and it can accurately capture difficult-to-model effects such as defocus, chromatic aberration, and higher-order aberrations with relatively few parameters. Overall, the lens blur field diminishes the tradeoff between representational accuracy and ease of acquisition by leveraging the interpolation properties of MLPs [67, 68], enabling scalable modeling within the input dimensions while accurately capturing detailed optical lens effects.
@@
Our work provides new insights and enables new applications. For example we show that PSFs vary considerably between smartphone models, including the iPhone 12 Pro and iPhone 14 Pro; that dual-pixel auto-focus systems have PSFs that depart significantly from the analytical models used hitherto; and that lens blur fields are reliable and expressive enough to reveal subtle differences in the optical behavior of smartphone devices of the same make and model, yielding a device-specific “blur signature”. We also show that lens blur fields can be used as space-efficient computational proxies of device optics, allowing us to apply high-dimensional blurs to render images with device-specific PSFs (Fig. 8). Taken together, we provide new ways to both model and simulate the performance of common imaging systems, including smartphone cameras used by billions worldwide.
Figure 2: Overview of learning lens blur fields. (A) Our approach learns a blur field by solving a non-blind deconvolution problem between blur-free calibration patterns $\hat{I}$ and blurry images $I$ of the patterns. Blur-free images are computed by applying perspective projection, lens distortion, and radiometric compensation to the albedo $A$ at the scene plane (top left). (B) An MLP is used to represent the lens blur field, which parameterizes the PSF over sensor position $\mathbf{x}$, target distance $d_{i}$, focus distance $f_{i}$, and a displacement $\mathbf{u}$ from $\mathbf{x}$. We train the MLP on a dataset of blur-free and captured focal stack images of synthetic noise patterns [69] and optimize MLP parameters to minimize the error between predicted and captured blurry image patches. Choice of noise patterns are discussed in Supplement Sec. S-IV.
@@ II-A Lens Blur Field Representation
Consider an incoherent point light source $\mathbf{p}\in\mathbb{R}^{3}$ at distance $d$ from the front pupil plane of a lens (see Fig. 2). Let $\mathbf{\tilde{x}}\in\mathbb{R}^{2}$ be the perspective projection of point $\mathbf{p}$ onto the camera’s sensor plane and let $\mathbf{x}\in\mathbb{R}^{2}$ be the point’s actual projection after accounting for non-linear geometric distortions caused by the lens. Since a camera’s focus setting $f$ may affect the overall magnification of captured images [70] and since dual-pixel AF sensors induce disparity between left-looking and right-looking pixels, in general $\mathbf{\tilde{x}}$ and $\mathbf{x}$ depend on both $f$ and the pixel type $c$. In the following, we express the one-to-one relation between geometrically-distorted points on the sensor plane and point sources at distance $d$ using
@@ II-B Continuous Image Formation
$\hat{I}^{(c)}(\mathbf{x},d,f)~{}=~{}A^{(c)}(\mathbf{p})~{}E^{(c)}(\mathbf{p},f){\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0},}$ (2)
@@
As illustrated in Fig. 2, we optimize the parameters of the blur field MLP by minimizing a photometric loss between a set of captured blurry images of known 2D patterns and corresponding rendered images estimated using Eq. 3.
@@
The outer summation is performed over all pixel channels $c$, all sensor points $\mathbf{x}$, focus and distance settings $f_{i}$ and $d_{i}$, $1\leq i\leq N_{I}$, and for each of $N_{T}$ known patterns ($1\leq t\leq N_{T}$). The inner summation is an approximation of the integral formulation of Eq. 3 with a discrete sum over $\mathbf{u}$. We consider two instances of the training procedure depending on whether or not the distance $d_{i}$ is the same for all input images. In the former case $d$ is fixed, yielding a 5D blur field; otherwise we optimize for a 6D blur field.
@@
We train on raw captured images, with the summation over $\mathbf{x}$ in Sec. II-B taken only over pixels for which we have image measurements for a given channel. For example, for a sensor with a Bayer color filter array, we supervise the red, green, and blue channels only on coordinates $\mathbf{x}$ that are explicitly measured.
The MLP uses seven fully-connected layers of 512 channels each with ReLU activation functions. The final layer has a sigmoid activation to satisfy the non-negativity constraint. Unlike recent work employing MLPs, we do not apply positional encoding to the MLP’s input coordinates [73] as we find it unnecessary for accurate PSF acquisition. Blur fields are trained for between $4\times 10^{6}$ and $10\times 10^{6}$ iterations with the Adam optimizer ($\beta_{1}=0.5$, $\beta_{2}=0.999$) on a single NVIDIA A6000 using the tiny-cuda-nn framework [74]. Details on other hyperparameters (learning rate, exponential decay, etc.) are included in Supplement Sec. S-III. Ablations on MLP architecture are included in Supplement Sec. S-IV.
@@
During training, we set one batch element to comprise a 2D PSF obtained by first randomly sampling a focus setting $f_{i}$ and a point $\mathbf{x}$ at a pixel center. Then we sample a grid of points in the two dimensions of $\mathbf{u}$ within the patch around $\mathbf{x}$ and query the MLP at these points to retrieve a 2D PSF. Sec. II-B is computed using the output of the convolution of the 2D PSF with $\hat{I}$ within the area of the patch. In our experiments, we use a batch size of one and allow 2D PSFs up to $120\times 120$ samples in size. The size of a patch $\mathcal{P}$ is set to be 1.5 times the size of the 2D PSF along each dimension, and the convolution is performed within the valid region of the patch to avoid boundary artifacts.
@@ III Image Acquisition Procedure
We describe the procedure to acquire the blur-free and captured blurry images of the calibration patterns used to optimize the lens blur field. The steps of the procedure are outlined in Fig. 2 (A) and consist of first computing the blur-free calibration image by applying (1) perspective projection via homography, (2) lens distortion, and (3) radiometric compensation to the scene-plane albedo $A$.
@@ III-A Homography Estimation
The perspective projection of the scene-plane albedo $A(\mathbf{p})$ onto the sensor plane (i.e., $\mathbf{p}\rightarrow\mathbf{\tilde{x}}$—see Fig. 2) is computed via a homography $H(\mathbf{\tilde{x}},d_{i},c)$. To estimate the homography, we require a pattern that enables robust point localization with sub-pixel accuracy, even in the presence of large defocus. To this end, we use two dot-grid calibration patterns consisting of white circles on a black background and black circles on a white background (shown, right). We compute the albedo dot-grid images and capture in-focus dot-grid images for which we also remove radial distortion using the procedure described in the next subsection. Following Scharstein and Szeliski [76], we binarize the dot-grid images and detect the pixel coordinates of the dot centers in each image. Then we refine the dot center locations in both images using a center-of-mass calculation on the dots in the original albedo and captured images. Finally, we compute the homography using RANSAC [77] between the dot center point correspondences in the images. Applying the homography to the scene-plane albedo as $A(H^{-1}(\mathbf{\tilde{x}},d_{i},c))$ gives the sensor-plane albedo.
@@ III-B Lens Distortion Estimation
To estimate the lens distortion we use the sixth-order Brown-Conrady model of radial distortion implemented in OpenCV’s calibrateCamera function [78] and apply it to the dot-grid center positions of the captured in-focus image described in the homography estimation step. We apply this model of radial distortion $D(\mathbf{\tilde{x}},c)$ to all scaled sensor-plane albedo images over $f_{i}$ and $d_{i}$. The composition of homography $H$, scale transformation $S$, and radial distortion $D$ define the invproj operator introduced in Eq. 1, which maps points $\mathbf{x}$ on the sensor plane to points $\mathbf{p}$ on the scene plane.
@@ III-C Radiometric Compensation
By combining Eq. 4 and its analogue for $I_{\text{max}}$ which follows trivially, the blur-free image $\hat{I}$ can now be expressed as an affine combination of $I_{\text{min}}$ and $I_{\text{max}}$.
@@ III-D Image Capture
Finally, we describe the procedure for capturing focal stack images displayed on a monitor. We use a 32-inch 4K LED monitor to display the calibration patterns for both smartphone and SLR captures. We mount each camera on a tripod and position it fronto-parallel to the monitor. Depending on the distance $d$ to the scene plane, the image of the calibration pattern may span the entire sensor plane or just a portion of it. In the former case, a single focal stack of images is captured for lens distortion estimation and radiometric compensation. In the latter, the camera and/or monitor are physically moved to maintain $d$ while also adjusting the patterns to project to different parts of the sensor plane (see result in Sec. IV). For smartphone experiments, focal stacks are automatically captured using software, and we sample a range of focus positions denoted as $N_{I}\in\{7,24\}$, which depends on hardware and device limitations. Finally, we set the ISO to the lowest possible value and capture bursts with exposure times of 100 ms (for iPhones) or 250 ms (for Pixel and SLR) to avoid capturing screen refresh, PWM, and flicker. Supplement Sec. S-II contains additional details on the acquisition procedure for all devices in our dataset.
Figure 3: Simulated recovery of spatially invariant dual-pixel blur kernels created with the parametric models proposed by Punnappurath et al. [10]. Each sub-panel gives a side-by-side comparison of the ground truth (left), our recovered PSF (middle left), the result of Joshi et al. [31] (middle right), and Mannan & Langer [32] (right). Root-mean-square error (RMSE) is estimated on the inset regions shown in Fig. 3 to mitigate bias from large regions of zero-valued pixels. The values are shown in the bottom left corner of each plot. Supplement Sec. S-IV contains more extensive comparisons and cross-sectional plots.